# Pricing energy spread options with variance gamma-driven Ornstein-Uhlenbeck dynamics

### v2

@@ 3 Analytical formula for cgf of the innovation term
Our main result is the analytical formula for the cgf of the innovation term ${\bf V}^{*}(t)$, $t>0$, of a multivariate OU-VG process.
This formula, which is stated in Theorem 3.3 and will later be applied under both $\mathbb{P}$ and $\mathbb{Q}_{{\bf h}}$, has 3 applications in the pricing of energy derivatives. Firstly, in the univariate and bivariate setting under $\mathbb{P}$, it is used in the likelihood function to estimate the LDOUP parameters (see Section 5.3). Secondly, in the univariate setting under $\mathbb{Q}_{{\bf h}}$, it is used to price forwards (see Section 4.4.1), calls (see Section 4.4.2), and calibrate the model (see Section 5.4). Thirdly, in the bivariate setting under $\mathbb{Q}_{{\bf h}}$, it is used to price spread options (see Sections 4.4.3 and 5.5).
@@
The random vector ${\bf Z}^{*}(t)$ is infinitely divisible, and by using the Lévy measure of $e^{-\lambda t}\widetilde{\bf Z}^{*}(t)$, where $\widetilde{\bf Z}={\bf Z}\circ(\lambda I)$, given by (sato99, , Lemma 17.1), it follows that the Lévy measure of ${\bf Z}^{*}(t)$ is
${\cal Z}^{*}(B)=\int_{\mathbb{R}^{n}\backslash\{{\bf 0}\}}\int_{0}^{t}\lambda{\bf 1}_{e^{\lambda t}B}(e^{-\lambda s}{\bf z})\,{\rm d}s{\cal Z}({\rm d}{\bf z}),\quad B\in{\cal B}(\mathbb{R}^{n}_{*}),$
@@
$I:=\int_{\mathbb{D}^{C}}e^{\langle\boldsymbol{\theta},{\bf z}\rangle}\,{\cal Z}^{*}({\rm d}{\bf z})$ $=\int_{\mathbb{D}^{C}}\int_{0}^{\lambda t}e^{\langle\boldsymbol{\theta},e^{u}{\bf z}\rangle}\,{\rm d}u{\cal Z}({\rm d}{\bf z})$
@@
for all $u\in[0,\lambda t]$. Next, by a version of the dominated convergence theorem (see (Ba01, , Lemma 16.1)) using ${\bf z}\mapsto 1+e^{\langle e^{\lambda t}\boldsymbol{\theta},{\bf z}\rangle}$ as the dominating function and the definition of the Lévy measure for $\cal Z$}, it follows that $f$ is continuous on $0,\lambda T]$. Thus, $I<\infty$, so that $\kappa_{{\bf Z}^{*}(t)}(\boldsymbol{\theta})$ is finite for all $\boldsymbol{\theta}\in{\cal D}_{{\bf Z}^{*}(t)}$ by the exponential moment theorem.
@@
Let $\boldsymbol{\theta}\in{\cal D}_{{\bf V}^{*}(t)}$ defined in Lemma 3.1. Consider the $\|\boldsymbol{\theta}\|_{{\boldsymbol{\Sigma}}}^{2}>0$ case. Let $A:=A(\boldsymbol{\theta})$, $B:=B(\boldsymbol{\theta})$, $c:=e^{\lambda t}$. Notice that
@@ 4.1 Energy market model
Let $\boldsymbol{\Lambda}=(\boldsymbol{\Lambda}(t))_{t\in[0,T]}$ be a deterministic seasonality function with $k$th marginal component
$k=1,\dots,n$, where $p$ is the period, and $b_{0k},b_{1k},b_{2k},b_{3k}\in\mathbb{R}$. Let
@@
with initial value ${\bf X}_{0}$ in its stationary distribution. Thus, the BDLP is ${\bf Z}\sim VGC^{n}_{m}({\bf b},\boldsymbol{\mu},{\boldsymbol{\Sigma}},\boldsymbol{\eta})$, and we work within this more general setting throughout, which includes ${\bf Z}\sim WVAG^{n}(a,\boldsymbol{\alpha},\allowbreak\boldsymbol{\mu},{\boldsymbol{\Sigma}},\boldsymbol{\eta})$ as a subclass. In Section 5.1, we make an additional assumption on $\boldsymbol{\eta}$, which is not needed now.
@@
We use a geometric model which ensures prices are positive. While electricity often exhibits negative intraday prices, derivative pricing is usually based on average daily price data (for example, see (bs14, , Section 3), (bsb04, , Section 3), (benthbook, , Section 9.3.2)), which are almost always positive.
@@ 4.2 Esscher transform
We now use the Esscher transform to obtain an EMM. For references on Esscher transforms and its applications to derivative pricing in equity markets see gs94 ; hub , (pasc, , Section 13.5), (shir, , Section VII.3c), and for applications to energy markets, see benthbook .
Let ${\bf Z}\sim L^{n}$ be a Lévy process. The Esscher measure $\mathbb{Q}_{{\bf h}}$ of $\bf Z$} with Esscher parameter or market price of risk (MPR) ${\bf h}=(h_{1},\dots,h_{n})$ is defined by the Radon-Nikodym derivative
@@
The no arbitrage condition that the discounted traded asset prices are $\mathbb{Q}_{{\bf h}}$-martingales trivially holds for the risk-free asset, which is the only traded asset, so that the Esssher transform provides up to infinitely many EMMs $\mathbb{Q}_{{\bf h}}$ indexed by ${\bf h}\in{\cal D}_{\bf Z}$, all consistent with no arbitrage. The market chooses a MPR ${\bf h}\in{\cal D}_{{\bf Z}}$ based on risk preferences, which we infer by calibrating the model to observed forward or call prices.
We now give some results on the $\mathbb{Q}_{{\bf h}}$-dynamics of the BDLP $\bf Z$}, which by equivalence determines the $\mathbb{Q}_{{\bf h}}$-dynamics of $\bf S$}. We begin with the case where the BDLP is multivariate VG. While the univariate case can be found in (hub, , Proposition 4), the multivariate case does not appear to be readily available except in (ant, , Theorem 6.3.1), and we reproduce this proof below. However, it is also an implication of (BKMS16, , Theorem 2.16).
@@
Suppose ${\bf V}\sim VG^{n}(b,\boldsymbol{\mu},{\boldsymbol{\Sigma}})$ under $\mathbb{P}$. If ${\bf h}\in{\cal D}_{{\bf V}}$, then ${\bf V}\sim VG^{n}(b_{h},\allowbreak\boldsymbol{\mu}_{\bf h},\Sigma_{\bf h})$ under $\mathbb{Q}_{\bf h}$, where
@@
Note that ${\cal D}_{{\bf V}}$ is defined in (2.3) using the parameters $(b,\boldsymbol{\mu},\Sigma)$ under $\mathbb{P}$. This is always the case, including similarly for the proceeding results.
We see that the existence of the Esscher measure $\mathbb{Q}_{{\bf h}}$ depends importantly on $\bf h$} and the kurtosis parameter $b>0$. For any choice of $(b,\boldsymbol{\mu},\Sigma)$, if $\|{\bf h}\|$ is sufficiently small, then $K_{\bf h}>0$, which is equivalent to ${\bf h}\in{\cal D}_{{\bf V}}$ and $\mathbb{Q}_{{\bf h}}$ existing. Alternatively, for any choice of $\bf h$}, if $b$ is sufficiently large, so that the kurtosis is sufficiently small, then $K_{\bf h}>0$. Otherwise, for an arbitrary choice of $(b,\boldsymbol{\mu},\Sigma)$ and $\bf h$}, $\mathbb{Q}_{{\bf h}}$ may not exist.
Let $p_{\bf h}$ denote the map sending the $\mathbb{P}$ parameter to the $\mathbb{Q}_{\bf h}$ parameters for the VG process, so $p_{\bf h}(b,\mu,{\boldsymbol{\Sigma}})=(b_{\bf h},\boldsymbol{\mu}_{\bf h},{\boldsymbol{\Sigma}}_{\bf h}).$
@@
${\bf Z}\stackrel{{\scriptstyle D}}{{=}}\boldsymbol{\eta}I+{\bf V}_{1}+\dots+{\bf V}_{m}.$
from (2.4). If ${\bf h}\in{\cal D}_{{\bf Z}}$, then this holds under $\mathbb{Q}_{\bf h}$ with ${\bf Z}\sim VGC^{n}_{m}({\bf b}_{\bf h},\boldsymbol{\mu}_{\bf h},{\boldsymbol{\Sigma}}_{\bf h},\boldsymbol{\eta})$, where ${\bf V}_{j}\sim VG^{n}([{\bf b}_{\bf h}]_{j},[\boldsymbol{\mu}_{\bf h}]_{j},[\Sigma_{\bf h}]_{j})$, $j=1,\dots,m$, are independent, and
$([{\bf b}_{\bf h}]_{j},[\boldsymbol{\mu}_{\bf h}]_{j},[\Sigma_{\bf h}]_{j})=p_{\bf h}(b_{j},\boldsymbol{\mu}_{j},\Sigma_{j}),$
@@
${\bf Z}\stackrel{{\scriptstyle D}}{{=}}\boldsymbol{\eta}I+{\bf V}_{0}+(V_{1},\dots,V_{n})$
@@
${[{\bf V}_{0}]}_{k}\sim VG^{1}(a,{[\boldsymbol{\mu}_{\bf h}]}_{k},{[\Sigma_{\bf h}]}_{kk}),\quad V_{k}\sim VG^{1}(\beta_{k},\mu_{h_{k}},\Sigma_{h_{k}}),$ (4.8)
@@ 4.3 Analytical formula for cgf of the log return
for a random vector ${\bf Z}^{*}\stackrel{{\scriptstyle D}}{{=}}{\bf Z}^{*}(\tau)$ independent of ${\cal F}_{t}$. Let $\kappa_{{\bf R}(t,T)}^{\mathbb{Q}_{\bf h}}(\boldsymbol{\theta}):=\kappa_{{\bf R}(t,T){\,|\,}{\cal F}_{t}}^{\mathbb{Q}_{\bf h}}(\boldsymbol{\theta})$ denote the cgf of this conditional distribution.
@@
for $\boldsymbol{\theta}\in{\cal D}_{{\bf Z}}$, where
$\kappa_{{\bf Z}^{*}(\tau)}^{\mathbb{Q}_{\bf h}}(\boldsymbol{\theta})={\rm i}\langle(e^{\lambda\tau}-1)\boldsymbol{\eta},\boldsymbol{\theta}\rangle+\sum_{j=1}^{m}\kappa_{{\bf V}_{j}^{*}(\tau)}^{\mathbb{Q}_{\bf h}}(\boldsymbol{\theta}),$ (4.11)
and the law of ${\bf V}_{j}$, $j=1,\dots,m$, under $\mathbb{Q}_{{\bf h}}$ is given in Corollary 4.2, and $\kappa_{{\bf V}_{j}^{*}(\tau)}^{\mathbb{Q}_{\bf h}}$ is evaluated by (3.3). Furthermore, (4.10) holds for $\boldsymbol{\theta}\in\mathbb{C}^{n}$ such that $\operatorname{Re}(\boldsymbol{\theta})\in{\cal D}_{{\bf Z}}$.
@@
which is finite by the assumption that $\boldsymbol{\theta}\in{\cal D}_{{\bf Z}}$ which is equivalent to $e^{-\lambda\tau}\boldsymbol{\theta}\in{\cal D}_{{\bf Z}^{*}(\tau)}$. Thus, we obtain (4.10). Next, (4.11) follows from Corollary 4.2, which gives the law of $\bf Z$} under $\mathbb{Q}_{{\bf h}}$, and (2.13) which gives the cgf of the corresponding $Z^{*}(\tau)$. Lastly, $\kappa_{{\bf R}(t,T)}^{\mathbb{Q}_{\bf h}}$ can be analytically continued by the exponential moment theorem for all $\boldsymbol{\theta}\in\mathbb{C}^{n}$ such that $\operatorname{Re}(\boldsymbol{\theta})\in{\cal D}_{{\bf Z}}$. ∎
@@
Using the setting and notation of Corollary 4.3, suppose ${\bf Z}\sim WVAG^{n}(a,\boldsymbol{\alpha},\boldsymbol{\mu},{\boldsymbol{\Sigma}},\boldsymbol{\eta})$ under $\mathbb{P}$ and ${\bf h}\in{\cal D}_{{\bf Z}}$. Then $\kappa_{{\bf R}(t,T)}^{\mathbb{Q}_{\bf h}}(\boldsymbol{\theta})$ follows (4.10) for $\boldsymbol{\theta}=(\theta_{1},\dots,\theta_{n})\in{\cal D}_{{\bf Z}}$, where
$\kappa_{{\bf Z}^{*}(\tau)}^{\mathbb{Q}_{\bf h}}(\boldsymbol{\theta})={\rm i}\langle(e^{\lambda\tau}-1)\boldsymbol{\eta},\boldsymbol{\theta}\rangle+\kappa_{{\bf V}_{0}^{*}(\tau)}^{\mathbb{Q}_{\bf h}}(\boldsymbol{\theta})+\sum_{i=1}^{n}\kappa_{V_{k}^{*}(\tau)}^{\mathbb{Q}_{\bf h}}(\theta_{k})$
where the law of ${\bf V}_{0}$, $V_{k}$, $k=1,\dots,n$, under $\mathbb{Q}_{{\bf h}}$ as given in Corollary 4.3, and $\kappa_{{\bf V}_{0}^{*}(\tau)}^{\mathbb{Q}_{\bf h}}$, $\kappa_{V_{k}^{*}(\tau)}^{\mathbb{Q}_{\bf h}}$ are evaluated by (3.3). Furthermore, (4.10) holds for $\boldsymbol{\theta}\in\mathbb{C}^{n}$ such that $\operatorname{Re}(\boldsymbol{\theta})\in{\cal D}_{{\bf Z}}$.
@@
Using the setting and notation of Corollary 4.2, suppose ${\bf Z}\sim VGC^{n}_{m}({\bf b},\boldsymbol{\mu},{\boldsymbol{\Sigma}},\boldsymbol{\eta})$ under $\mathbb{P}$ and ${\bf h}\in{\cal D}_{{\bf Z}}$. Then $V_{jk}:={[{\bf V}_{0}]}_{k}\sim VG^{1}({[[{\bf b}_{\bf h}]_{j}]}_{k},\allowbreak{[[\boldsymbol{\mu}_{\bf h}]_{j}]}_{k},{[[{\boldsymbol{\Sigma}}_{\bf h}]_{j}]}_{kk})$, $j=1,\dots,m$, and
$\kappa^{\mathbb{Q}_{\bf h}}_{R_{k}(t,T)}(\theta)=(\Lambda(T)+X_{k}(t)e^{-\lambda\tau}-Y_{k}(t))\theta+\kappa^{\mathbb{Q}_{\bf h}}_{Z_{k}^{*}(\tau)}(e^{-\lambda\tau}\theta)$ (4.12)
for $\theta\in\mathbb{R}$ such that $\theta{\bf e}_{k}\in{\cal D}_{\bf Z}$, where
@@
and $\kappa^{\mathbb{Q}_{\bf h}}_{V_{jk}^{*}(\tau)}$ is evaluated by (3.7). Furthermore, (4.12) holds for $\theta\in\mathbb{C}$ such that $\operatorname{Re}(\theta){\bf e}_{k}\in{\cal D}_{{\bf Z}}$.
@@
Using the setting and notation of Corollary 4.3, suppose ${\bf Z}\sim WVAG^{n}(a,\boldsymbol{\alpha},\boldsymbol{\mu},{\boldsymbol{\Sigma}},\boldsymbol{\eta})$ under $\mathbb{P}$ and ${\bf h}\in{\cal D}_{{\bf Z}}$. Then $V_{0k}:={[{\bf V}_{0}]}_{k}$ and $V_{k}$ satisfies (4.8), and $\kappa^{\mathbb{Q}_{\bf h}}_{R_{k}(t,T)}(\theta)$ follows (4.12) for $\theta\in\mathbb{R}$ such that $\theta{\bf e}_{k}\in{\cal D}_{\bf Z}$, where
@@
and $\kappa^{\mathbb{Q}_{\bf h}}_{V_{0k}^{*}(\tau)}$, $\kappa^{\mathbb{Q}_{\bf h}}_{V_{k}^{*}(\tau)}$ are evaluated by (3.7). Furthermore, (4.12) holds for $\theta\in\mathbb{C}$ such that $\operatorname{Re}(\theta){\bf e}_{k}\in{\cal D}_{{\bf Z}}$.

### v3

@@ 3 Analytical formula for cgf of the innovation term
Our main result is the analytical formula for the cgf of ${\bf V}^{*}(t)$, $t>0$, the innovation term of a multivariate OU-VG process.
This formula, which is stated in Theorem 3.3 and will later be applied under both $\mathbb{P}$ and $\mathbb{Q}_{{\bf h}}$, has 3 applications in the pricing of energy derivatives. Firstly, in the univariate and bivariate setting under $\mathbb{P}$, it is used in the likelihood function to estimate the LDOUP parameters (see Section 5.1.3). Secondly, in the univariate setting under $\mathbb{Q}_{{\bf h}}$, it is used to price forwards (see Section 4.4.1), calls (see Section 4.4.2), and calibrate the model (see Section 5.1.4). Thirdly, in the bivariate setting under $\mathbb{Q}_{{\bf h}}$, it is used to price spread options (see Sections 4.4.3 and 5.1.5).
@@
[…]
${\cal Z}^{*}(B)=\int_{\mathbb{R}^{n}\backslash\{{\bf 0}\}}\int_{0}^{t}\lambda{\bf 1}_{e^{\lambda t}B}(e^{-\lambda s}{\bf z})\,{\rm d}s{\cal Z}({\rm d}{\bf z})$
@@
$J:=\int_{\mathbb{D}^{C}}e^{\langle\boldsymbol{\theta},{\bf z}\rangle}\,{\cal Z}^{*}({\rm d}{\bf z})$ $=\int_{\mathbb{D}^{C}}\int_{0}^{\lambda t}e^{\langle\boldsymbol{\theta},e^{u}{\bf z}\rangle}\,{\rm d}u{\cal Z}({\rm d}{\bf z})$
@@
for all $u\in[0,\lambda t]$. Next, by a version of the dominated convergence theorem (see (Ba01, , Lemma 16.1)) using ${\bf z}\mapsto 1+e^{\langle e^{\lambda t}\boldsymbol{\theta},{\bf z}\rangle}$ as the dominating function and the definition of the Lévy measure for $\cal Z$}, it follows that $f$ is continuous on $0,\lambda T]$. Thus, $J<\infty$, so that $\kappa_{{\bf Z}^{*}(t)}(\boldsymbol{\theta})$ is finite for all $\boldsymbol{\theta}\in{\cal D}_{{\bf Z}^{*}(t)}$ by the exponential moment theorem.
@@
Let $\boldsymbol{\theta}\in{\cal D}_{{\bf V}^{*}(t)}$. Consider the $\|\boldsymbol{\theta}\|_{{\boldsymbol{\Sigma}}}^{2}>0$ case. Let $A:=A(\boldsymbol{\theta})$, $B:=B(\boldsymbol{\theta})$, $c:=e^{\lambda t}$. Notice that
@@ 4.1 Energy market model
Let $\boldsymbol{\Lambda}=(\boldsymbol{\Lambda}(t))_{t\in[0,T]}=(\Lambda_{1},\dots,\Lambda_{n})$ be a deterministic seasonality function. Let
@@
with initial value ${\bf X}_{0}$ in its stationary distribution. Thus, the BDLP is ${\bf Z}\sim VGC^{n}_{m}({\bf b},\boldsymbol{\mu},{\boldsymbol{\Sigma}},\boldsymbol{\eta})$, and we work within this more general setting throughout, which includes ${\bf Z}\sim WVAG^{n}(a,\boldsymbol{\alpha},\allowbreak\boldsymbol{\mu},{\boldsymbol{\Sigma}},\boldsymbol{\eta})$ as a subclass. In Section 5.1.1, we make an additional assumption on $\boldsymbol{\eta}$, which is not needed now.
@@
This is known as a geometric model and ensures that prices are positive. In contrast, an arithmetic model would replace (4.1) with ${\bf S}(t)=\boldsymbol{\Lambda}(t)+{\bf X}(t)$. While electricity often exhibits negative intraday prices, derivative pricing is usually based on average daily price data (for example, see (bs14, , Section 3), (bsb04, , Section 3), (benthbook, , Section 9.3.2)), which are almost always positive. However, in Section 5.2, we have a real data example with negative prices and discuss a trick that allows a geometric model to still be used.
@@ 4.2 Esscher transform
We now use the Esscher transform to obtain an EMM. For references on Esscher transforms and its applications to derivative pricing in equity markets, see gs94 ; hub , (pasc, , Section 13.5), (shir, , Section VII.3c), and for applications to energy markets, see benthbook .
Let ${\bf Z}\sim L^{n}$ be a Lévy process. The Esscher measure $\mathbb{Q}_{{\bf h}}$ of $\bf Z$} with market price of risk (MPR) (or Esscher parameter) ${\bf h}=(h_{1},\dots,h_{n})$ is defined by the Radon-Nikodym derivative
@@
The no arbitrage condition that the discounted traded asset prices are $\mathbb{Q}_{{\bf h}}$-martingales trivially holds for the risk-free asset, which as we recall is the only traded asset, so that the Esssher transform provides up to infinitely many EMMs $\mathbb{Q}_{{\bf h}}$ indexed by ${\bf h}\in{\cal D}_{\bf Z}$, all consistent with no arbitrage. The market chooses a MPR ${\bf h}\in{\cal D}_{{\bf Z}}$ based on risk preferences, which we infer by calibrating the model to observed forward or call prices.
We now give some results on the $\mathbb{Q}_{{\bf h}}$-dynamics of the BDLP $\bf Z$}, which by equivalence determines the $\mathbb{Q}_{{\bf h}}$-dynamics of $\bf S$}. We begin by showing how the parameters of the multivariate VG process change under the Esscher transform. While the univariate case can be found in (hub, , Proposition 4), the multivariate case does not appear to be readily available except in (ant, , Theorem 6.3.1), and we reproduce this proof below. However, it is also an implication of (BKMS16, , Theorem 2.16).
@@
Suppose ${\bf V}\sim VG^{n}(b,\boldsymbol{\mu},{\boldsymbol{\Sigma}})$ under $\mathbb{P}$. If ${\bf h}\in{\cal D}_{{\bf V}}$, then ${\bf V}\sim VG^{n}(b_{h},\allowbreak\boldsymbol{\mu}_{\bf h},{\boldsymbol{\Sigma}}_{\bf h})$ under $\mathbb{Q}_{\bf h}$, where
@@
Note that ${\cal D}_{{\bf V}}$ is defined in (2.3) using the parameters $(b,\boldsymbol{\mu},{\boldsymbol{\Sigma}})$ under $\mathbb{P}$. This is always the case, including similarly for the proceeding results.
We see that the existence of the Esscher measure $\mathbb{Q}_{{\bf h}}$ depends importantly on $\bf h$} and the kurtosis parameter $b>0$. For any choice of $(b,\boldsymbol{\mu},{\boldsymbol{\Sigma}})$, if $\|{\bf h}\|$ is sufficiently small, then $K_{\bf h}>0$, which is equivalent to ${\bf h}\in{\cal D}_{{\bf V}}$ and $\mathbb{Q}_{{\bf h}}$ existing. Alternatively, for any choice of $\bf h$}, if $b$ is sufficiently large, so that the kurtosis is sufficiently small, then $K_{\bf h}>0$. Otherwise, for an arbitrary choice of $(b,\boldsymbol{\mu},{\boldsymbol{\Sigma}})$ and $\bf h$}, $\mathbb{Q}_{{\bf h}}$ may not exist.
Let $p_{\bf h}$ denote the map sending the $\mathbb{P}$ parameter to the $\mathbb{Q}_{\bf h}$ parameters for the VG process, so $p_{\bf h}(b,\mu,{\boldsymbol{\Sigma}}):=(b_{\bf h},\boldsymbol{\mu}_{\bf h},{\boldsymbol{\Sigma}}_{\bf h}).$
@@
${\bf Z}(t)=\boldsymbol{\eta}t+{\bf V}_{1}(t)+\dots+{\bf V}_{m}(t),\quad t\geq 0.$
from (2.4). If ${\bf h}\in{\cal D}_{{\bf Z}}$, then this holds under $\mathbb{Q}_{\bf h}$ with ${\bf Z}\sim VGC^{n}_{m}({\bf b}_{\bf h},\boldsymbol{\mu}_{\bf h},{\boldsymbol{\Sigma}}_{\bf h},\boldsymbol{\eta})$, where ${\bf V}_{j}\sim VG^{n}([{\bf b}_{\bf h}]_{j},[\boldsymbol{\mu}_{\bf h}]_{j},[{\boldsymbol{\Sigma}}_{\bf h}]_{j})$, $j=1,\dots,m$, are independent, and
$([{\bf b}_{\bf h}]_{j},[\boldsymbol{\mu}_{\bf h}]_{j},[{\boldsymbol{\Sigma}}_{\bf h}]_{j})=p_{\bf h}(b_{j},\boldsymbol{\mu}_{j},{\boldsymbol{\Sigma}}_{j}),$
@@
${\bf Z}(t)=\boldsymbol{\eta}t+{\bf V}_{0}(t)+(V_{1}(t),\dots,V_{n}(t)),\quad t\geq 0,$
@@
${[{\bf V}_{0}]}_{k}\sim VG^{1}(a,{[\boldsymbol{\mu}_{\bf h}]}_{k},{[{\boldsymbol{\Sigma}}_{\bf h}]}_{kk}),\quad V_{k}\sim VG^{1}(\beta_{k},\mu_{h_{k}},\Sigma_{h_{k}}),$ (4.8)
@@ 4.3 Analytical formula for cgf of the log return
for a random vector ${\bf Z}^{*}\stackrel{{\scriptstyle D}}{{=}}{\bf Z}^{*}(\tau)$ independent of ${\cal F}_{t}$. Let $\kappa_{{\bf R}(t,T)}^{\mathbb{Q}_{\bf h}}(\boldsymbol{\theta}):=\linebreak\kappa_{{\bf R}(t,T){\,|\,}{\cal F}_{t}}^{\mathbb{Q}_{\bf h}}(\boldsymbol{\theta})$ denote the cgf of this conditional distribution.
For a Lévy process ${\bf Z}\sim L^{n}$, recall the notation ${\cal D}_{{\bf Z}}^{\mathbb{Q}_{{\bf h}}}:=\{\boldsymbol{\theta}\in\mathbb{R}^{n}:\linebreak\mathbb{E}_{\mathbb{Q}_{{\bf h}}}[e^{\langle\boldsymbol{\theta},{\bf Z}(1)\rangle}]<\infty\}$. It follows from (pasc, , Theorem 13.66) that $\boldsymbol{\theta}\in{\cal D}_{{\bf Z}}^{\mathbb{Q}_{{\bf h}}}$ is equivalent to $\boldsymbol{\theta}+{\bf h}\in{\cal D}_{{\bf Z}}$, which expresses the domain of $\kappa_{{\bf Z}}^{\mathbb{Q}_{\bf h}}$ under $\mathbb{Q}_{\bf h}$ in terms of the domain ${\cal D}_{{\bf Z}}$ under $\mathbb{P}$, which is the approach we mainly use. Alternatively, from Corollary 4.2, noting that ${\bf Z}\sim VGC^{n}_{m}$ is closed under Esscher transform but with different parameters under $\mathbb{P}$ and $\mathbb{Q}_{{\bf h}}$, ${\cal D}_{{\bf Z}}^{\mathbb{Q}_{{\bf h}}}$ can be understood as ${\cal D}_{{\bf Z}}$ but with the $\mathbb{P}$-parameters replaced by the $\mathbb{Q}_{{\bf h}}$-parameters.
@@
for $\boldsymbol{\theta}+{\bf h}\in{\cal D}_{{\bf Z}}$, where
$\kappa_{{\bf Z}^{*}(\tau)}^{\mathbb{Q}_{\bf h}}(\boldsymbol{\theta})=\langle(e^{\lambda\tau}-1)\boldsymbol{\eta},\boldsymbol{\theta}\rangle+\sum_{j=1}^{m}\kappa_{{\bf V}_{j}^{*}(\tau)}^{\mathbb{Q}_{\bf h}}(\boldsymbol{\theta}),$ (4.11)
and the law of ${\bf V}_{j}$, $j=1,\dots,m$, under $\mathbb{Q}_{{\bf h}}$ is given in Corollary 4.2, and $\kappa_{{\bf V}_{j}^{*}(\tau)}^{\mathbb{Q}_{\bf h}}$ is evaluated by (3.3). Furthermore, (4.10) holds for $\boldsymbol{\theta}\in\mathbb{C}^{n}$ such that $\operatorname{Re}(\boldsymbol{\theta})+{\bf h}\in{\cal D}_{{\bf Z}}$.
@@
which is finite since the assumption $\boldsymbol{\theta}+{\bf h}\in{\cal D}_{{\bf Z}}$ is equivalent to $\boldsymbol{\theta}\in{\cal D}_{{\bf Z}}^{\mathbb{Q}_{\bf h}}$, which is equivalent to $e^{-\lambda\tau}\boldsymbol{\theta}\in{\cal D}_{{\bf Z}^{*}(\tau)}^{\mathbb{Q}_{\bf h}}$. Thus, we obtain (4.10). Next, (4.11) follows from Corollary 4.2, which gives the law of $\bf Z$} under $\mathbb{Q}_{{\bf h}}$, and (2.12) which gives the cgf of the corresponding ${\bf Z}^{*}(\tau)$. Lastly, $\kappa_{{\bf R}(t,T)}^{\mathbb{Q}_{\bf h}}$ can be analytically continued by the exponential moment theorem for all $\boldsymbol{\theta}\in\mathbb{C}^{n}$ such that $\operatorname{Re}(\boldsymbol{\theta})+{\bf h}\in{\cal D}_{{\bf Z}}$. ∎
@@
Using the setting and notation of Corollary 4.3, suppose ${\bf Z}\sim WVAG^{n}(a,\boldsymbol{\alpha},\boldsymbol{\mu},{\boldsymbol{\Sigma}},\boldsymbol{\eta})$ under $\mathbb{P}$ and ${\bf h}\in{\cal D}_{{\bf Z}}$. Then $\kappa_{{\bf R}(t,T)}^{\mathbb{Q}_{\bf h}}(\boldsymbol{\theta})$ follows (4.10) for $\boldsymbol{\theta}+{\bf h}\in{\cal D}_{{\bf Z}}$, where
$\kappa_{{\bf Z}^{*}(\tau)}^{\mathbb{Q}_{\bf h}}(\boldsymbol{\theta})=\langle(e^{\lambda\tau}-1)\boldsymbol{\eta},\boldsymbol{\theta}\rangle+\kappa_{{\bf V}_{0}^{*}(\tau)}^{\mathbb{Q}_{\bf h}}(\boldsymbol{\theta})+\sum_{i=1}^{n}\kappa_{V_{k}^{*}(\tau)}^{\mathbb{Q}_{\bf h}}(\theta_{k})$
where the law of ${\bf V}_{0}$, $V_{k}$, $k=1,\dots,n$, under $\mathbb{Q}_{{\bf h}}$ as given in Corollary 4.3, and $\kappa_{{\bf V}_{0}^{*}(\tau)}^{\mathbb{Q}_{\bf h}}$, $\kappa_{V_{k}^{*}(\tau)}^{\mathbb{Q}_{\bf h}}$ are evaluated by (3.3). Furthermore, (4.10) holds for $\boldsymbol{\theta}\in\mathbb{C}^{n}$ such that $\operatorname{Re}(\boldsymbol{\theta})+{\bf h}\in{\cal D}_{{\bf Z}}$.
@@
Using the setting and notation of Corollary 4.2, suppose ${\bf Z}\sim VGC^{n}_{m}({\bf b},\boldsymbol{\mu},{\boldsymbol{\Sigma}},\boldsymbol{\eta})$ under $\mathbb{P}$ and ${\bf h}\in{\cal D}_{{\bf Z}}$. Then $V_{jk}:={[{\bf V}_{j}]}_{k}\sim VG^{1}({[[{\bf b}_{\bf h}]_{j}]}_{k},\allowbreak{[[\boldsymbol{\mu}_{\bf h}]_{j}]}_{k},{[[{\boldsymbol{\Sigma}}_{\bf h}]_{j}]}_{kk})$, $j=1,\dots,m$, and
$\kappa^{\mathbb{Q}_{\bf h}}_{R_{k}(t,T)}(\theta)=(\Lambda_{k}(T)+X_{k}(t)e^{-\lambda\tau}-Y_{k}(t))\theta+\kappa^{\mathbb{Q}_{\bf h}}_{Z_{k}^{*}(\tau)}(e^{-\lambda\tau}\theta)$ (4.12)
for $\theta\in\mathbb{R}$ such that $\theta{\bf e}_{k}+{\bf h}\in{\cal D}_{\bf Z}$, where
@@
and $\kappa^{\mathbb{Q}_{\bf h}}_{V_{jk}^{*}(\tau)}$ is evaluated by (3.7). Furthermore, (4.12) holds for $\theta\in\mathbb{C}$ such that $\operatorname{Re}(\theta){\bf e}_{k}+{\bf h}\in{\cal D}_{{\bf Z}}$.
@@
Using the setting and notation of Corollary 4.3, suppose ${\bf Z}\sim WVAG^{n}(a,\boldsymbol{\alpha},\boldsymbol{\mu},{\boldsymbol{\Sigma}},\boldsymbol{\eta})$ under $\mathbb{P}$ and ${\bf h}\in{\cal D}_{{\bf Z}}$. Then $V_{0k}:={[{\bf V}_{0}]}_{k}$ and $V_{k}$ satisfies (4.8), and $\kappa^{\mathbb{Q}_{\bf h}}_{R_{k}(t,T)}(\theta)$ follows (4.12) for $\theta\in\mathbb{R}$ such that $\theta{\bf e}_{k}+{\bf h}\in{\cal D}_{\bf Z}$, where
@@
and $\kappa^{\mathbb{Q}_{\bf h}}_{V_{0k}^{*}(\tau)}$, $\kappa^{\mathbb{Q}_{\bf h}}_{V_{k}^{*}(\tau)}$ are evaluated by (3.7). Furthermore, (4.12) holds for $\theta\in\mathbb{C}$ such that $\operatorname{Re}(\theta){\bf e}_{k}+{\bf h}\in{\cal D}_{{\bf Z}}$.